Please help. This one I think I have solved it, but I am unsure of what the solution means. Here is the problem.
Determine the nature of the solution. x^2-20x+100=0 a)2 non-real solutions
b)2 real solutions or c)1 real solution.
I know that if the answer is zero than it is 1 real, if it is positive, than it is 2 real, and if it is negative than it is 2 non-real. I think I got the solution down to 10+or- the square root of 0.
Is that right? Thank you for your help

Answer by stanbon(75887)

Determine the nature of the solution. x^2-20x+100=0
a)2 non-real solutions
b)2 real solutions or
c)1 real solution.
I know that if the answer is zero than it is 1 real, if it is positive, than it is 2 real, and if it is negative than it is 2 non-real. I think I got the solution down to 10+or- the square root of 0.
Is that right?
----------------
What you say is correct.
-----------------
b^2 - 4ac = 20^2 -4*1*100 = 400 - 400 = 0
So there is 1 Real Number answer.
By the way, that one answer has "multiplicity two".
=====================================
Cheers,
Stan H.